Show that the wavefunction Ψ(x) = N sin kx + iN cos kx is an eigen function of the momentum operator and determine its eigen value.


1
Expert's answer
2021-06-07T09:32:11-0400

The momentum operator is "\hat{p} =- i\hbar\dfrac{\partial}{\partial x}". By definition, the eigen function has the following property:


"\hat{p}\Psi(x) = \lambda\Psi(x)"

where "\lambda" is the corresponding eigenvalue. Substituting the function and peforming the differentiation, obtain:

"\hat{p}\Psi(x) = - i\hbar\dfrac{\partial}{\partial x} (N\sin(kx) + iN\cos(kx)) = - i\hbar(Nkcos(kx) - iNk\sin(kx)) =\\
=-i\hbar Nkcos(kx) - \hbar Nk\sin(kx)"

Factoring out "-\hbar k" , have:


"\hat{p}\Psi(x) = -\hbar k(N\sin(kx) + iN\cos(kx)) = -\hbar k\Psi(x)"


Thus, this finction is indeed the eigen function of the momentum operator. Comparing the last equation with the first one, find the eigenvalue:


"\lambda = -\hbar k"

Answer. "\Psi(x)" is the eigenfunction of the momentum operator with the eigenvalue "\lambda = -\hbar k".